a 3.0g mixture of na2co3 and calcium trioxocarbonate(iv) liberated 0.44g of co2 on strong heating until no gas was evolved.what is the percentage of na2co3
1
Expert's answer
2020-11-10T14:04:12-0500

Out of these two compounds only CaCO3 decomposes upon heating according to the reaction

CaCO3 --> CaO + CO2

Therefore,

"m_{CaCO_3}=\frac{0.44g CO_2}{44gCO_2/mol}\times\frac{1molCaCO_3}{1molCO_2}\times\frac{100.1gCaCO_3}{1molCaCO_3}=1.0g"


"\%CaCO_3=\frac{1.0g}{3.0g}\times100\%=33\%"


"\%Na_2CO_3=100-33=67\%"


Answer: 67%